# Are We in the AI-Generated Text World Already? Quantifying and Monitoring AIGT on Social Media

### v1

@@ Abstract
Social media platforms are experiencing a growing presence of AI-Generated Texts (AIGTs). However, the misuse of AIGTs could have profound implications for public opinion, such as spreading misinformation and manipulating narratives. Despite its importance, a systematic study to assess the prevalence of AIGTs on social media is still lacking. To address this gap, this paper aims to quantify, monitor, and analyze the AIGTs on online social media platforms. We first collect a dataset (SM-D) with around $2.4M$ posts from $3$ major social media platforms: Medium, Quora, and Reddit. Then, we construct a diverse dataset (AIGTBench) to train and evaluate AIGT detectors. AIGTBench combines popular open-source datasets and our AIGT datasets generated from social media texts by $12$ LLMs, serving as a benchmark for evaluating mainstream detectors. With this setup, we identify the best-performing detector (OSM-Det). We then apply OSM-Det to SM-D to track AIGTs over time and observe different trends of AI Attribution Rate (AAR) across social media platforms from January 2022 to October 2024. Specifically, Medium and Quora exhibit marked increases in AAR, rising from $1.77\%$ to $37.03\%$ and $2.06\%$ to $38.95\%$, respectively. In contrast, Reddit shows slower growth, with AAR increasing from $1.31\%$ to $2.45\%$ over the same period. Our further analysis indicates that AIGTs differ from human-written texts across several dimensions, including linguistic patterns, topic distributions, engagement levels, and the follower distribution of authors. We envision our analysis and findings on AIGTs in social media can shed light on future research in this domain.
@@ Introduction
Currently, numerous detectors have been developed to detect AIGTs. According to the MGTBench [14], these detectors are broadly divided into two categories: metric-based [9, 28] and model-based detectors [16, 39, 4], some of which have shown high accuracy and robustness. While these detectors have been applied in controlled settings, recent studies have explored their effectiveness in real-world scenarios. Hanley et al.[13] conduct AIGT detection on news website articles, with a primary focus on content generated by GPT-3.5 and others from Turing benchmark, which includes various pre-2022 models [46]. Furthermore, Liu et al.[24] carry out detection tests for ChatGPT-generated content on arXiv papers. However, they do not consider recent popular models, such as Llama [44] and GPT-4 [31], which also possess powerful text generation capabilities and are widely adopted. We thereby consider a broader range of models in our efforts to detect AIGTs on social media.
@@
To quantify and monitor AIGTs on social media, we collect textual data across $3$ popular platforms ranging from January 1, 2022, to October 31, 2024, as most LLMs are released after 2022. After data preprocessing, we obtain $1,170,821$ posts from Medium, $245,131$ answers from Quora, and $982,440$ comments from Reddit. We name it as SM-D, short for Social Media Dataset.
@@
To identify the most effective detector, we construct a dataset named AIGTBench, which consists of public AIGT/Supervised-Finetuning (SFT) datasets and our own AIGT datasets generated from social media data. AIGTBench includes AIGTs generated by $12$ different LLMs, such as GPT Series (GPT-3.5, GPT-4 and GPT4o-mini [32]), Llama Series (Llama-1, 2, 3 [44, 45, 7]), etc, totaling around $28.77M$ AIGT and $13.55M$ HWT samples. Building on AIGTBench, we benchmark AIGT detectors and leverage the best-performing detector as our primary detector for social media AIGT detection, which achieves an accuracy of $0.979$ and an F1-score of $0.980$ on AIGTBench. To better reflect its application in detecting content on online social media, we rename it as OSM-Det (Online Social Media Detector).
@@
Based on OSM-Det, we quantify and monitor the texts across the $3$ platforms and use the AI Attribution Rate (AAR) to represent the proportion of articles classified as AI-generated. We observe several noteworthy phenomena: (1) A sharp rise in AI-generated content begins in December 2022, with distinct AAR trends emerging across platforms. Before December 2022, the AAR across platforms remains stable. However, starting in December, Medium and Quora show significant surges, while Reddit shows only a slight increase. This suggests the widespread and diverse LLM adoption on social media; (2) Linguistic analysis shows similar AAR trends and exhibits stylistic features in AIGTs/HWTs. Based on the word-level analysis, we find that the usage trend of top-frequency AI-preferred words aligns closely with LLM adoption trends. With sentence-level analysis, we also reveal that AIGTs tend to be more objective and standardized, whereas HWTs are more flexible and informal; (3) Technology-related topics drive higher AARs on Medium. Topics like “Technology” and “Software Development” show the highest AARs, indicating that users with a strong technical background are more likely to adopt LLMs; (4) Predicted HWTs receive more engagement than AIGTs. On Medium, the content predicted by our OSM-Det as HWTs receives more average “Likes” and “Comments” than AIGTs. This suggests that users are more inclined to engage with HWTs; and (5) Authors with fewer followers are more likely to produce AIGTs. On Medium, users with no more than one thousand followers tend to produce content that has the highest mean AAR at $54.02\%$. In contrast, as the follower count increases, the AAR gradually shifts toward the lower range ($\leq 25.00\%$).
@@
The growth in model parameters and training data has recently empowered LLMs to demonstrate exceptional language processing capabilities and few-shot learning abilities [51]. Since then, LLMs have gradually gained popularity, like GPT-4 [31] and Llama [44], enabling users to generate high-quality texts effortlessly. Yet, LLMs have raised concerns about potential misuse, such as fake news generation [50], academic misconduct [47], and performance degradation of training LLMs using AI content [5], making the detection of AI-Generated texts (AIGTs, also known as machine-generated texts) increasingly important [8]. He et al.[14] introduce the first benchmark, MGTBench, for standardizing the evaluation of different LLMs and experimental setups within the AIGT detectors. They broadly categorize the detectors into two main types: metric-based and model-based detectors. Metric-based detectors use pre-defined metrics, such as log-likelihood values and rankings, to capture the characteristics of texts and identify AIGTs [9, 28, 41]. In contrast, model-based detectors rely on trained models to distinguish between AIGTs and HWTs [39, 12, 4, 10, 24, 16, 20]. For more introduction, refer to Appendix B.
Based on these AIGT detectors, some researchers have applied them to text detection in real-world scenarios. Hanley et al.[13] train a detector using data generated by the ChatGPT and Turing benchmark model and conduct detection tests on multiple news websites. Their study reveals that, from January 1, 2022, to May 1, 2023, the proportion of synthetic articles increased on news sites. Liu et al.[24] also conduct detection tests on arXiv and find a significant rise in the proportion of papers using ChatGPT-generated content, reaching $26.1\%$ by December 2023. In contrast to their detection targets, we focus on detecting AIGTs on social media platforms and covering a broader range of LLMs.
Macko et al.[25] construct a multilingual dataset based on instant messaging and social interaction platforms such as Telegram, Discord, and WhatsApp, using it to compare the performance of existing detectors. In contrast, our research focuses on providing an in-depth temporal analysis of AIGTs on content-driven social platforms like Medium, Quora, and Reddit.
@@
Table 1: Overview of the Medium, Quora, and Rediit datasets.
@@
Figure 1: Proportion of total sentences various LLMs, with “Others” including Alpaca 7B and Vicuna 13B.
To train the AIGT detectors, we consider two parts of the data. First, we consider $6$ publicly available AIGT datasets and $5$ common SFT datasets to form the training dataset (see Tables A1 and A2 for dataset statistics and Appendix D for more details). Second, to increase the detector’s generalization capabilities on social media, we additionally collect data from the $3$ social media platforms ranging from January 1, 2018, to December 31, 2021. We classify this data as HWTs, given that most LLMs had not been published during this period. We also design different LLMs writing tasks to generate AIGTs that align with the characteristics of platforms (Table A3 describes the statistics details).
@@
Overall, the datasets used for training our detector and the distribution of LLM series are shown in Figure 1. This dataset includes $12$ different LLMs, with a detailed introduction provided in Appendix A. Within these datasets, the two most prevalent model series are the GPT Series, which accounts for $42.99\%$, and the Llama series, which represents $39.05\%$. GPT Series is the most widely used proprietary model and has played a pivotal role in the evolution of generative AI. As of January 2023, approximately $13M$ users interact daily with GPT-3.5 [49]. The Llama series models also have significant influences, as the report indicates that downloads of Llama models on the Hugging Face platform have nearly reached around $350M$ [27]. Therefore, these two model series are the primary focus of our dataset. During the data generation process, we notice that certain samples contain textual noise, like irrelevant or redundant information. To maintain data quality, we implement some data processing strategies (see Appendix E for details).
Experimental Settings
Datasets
As mentioned in Section 3, we collect the social media dataset (SM-D) and the detector training dataset (AIGTBench). SM-D refers to the social media dataset that we conduct the quantification, with more details provided in Section 3.1. AIGTBench is the benchmark dataset for AIGT detectors, which includes samples generated by $12$ different LLMs, as described in Section 3.2. We randomly divide AIGTBench into training, validation, and test sets in a $7:1:2$ ratio. Specifically, the distribution of tokens across the texts in the training set is shown in Figure A1, and the validation and test sets maintain a consistent token distribution with the training set.
AIGT Detectors
During the detection process, we initially use the GPT-2 medium to extract multiple metrics, including log-likelihood and log-rank. Based on these extracted metrics, we train logistic regression models to enhance the accuracy of predictions.
For the model-based detectors, we consider both pre-trained detectors and fine-tuned models with the AIGTBench, that is, OpenAI Detector [39], ChatGPT Detector [12], ConDA [4], GPTZero [10], CheckGPT [24], and LM-D [16]. Specifically, for the OpenAI Detector and ChatGPT Detector, we consider their pre-trained version and select the RoBERTa-base model as it demonstrates stable performance across multiple detection tasks and typically provides better detection results. For ConDA and LM-D, we choose the Longformer-base-4096 model as the base model and fine-tune it with the AIGTBench. For GPTZero, we directly use its commercial API. For CheckGPT, we retrain the original training framework [24].
@@
Table 2: Performance of detectors on AIGTBench.
@@
To evaluate the performance of different detectors, we use accuracy and F1-score as the evaluation metrics, which are common standards in AIGT detection tasks. Besides, we introduce two new metrics AI Attribution Rate (AAR) and False Positive Rate (FPR) for social media text detection. The AAR indicates the proportion of texts that the model predicts as AI-generated, while the FPR measures the proportion of HWTs that are mistakenly identified as AIGTs.
To evaluate word usage, we calculate the term frequency and divide it by the total number of documents, obtaining the normalized term frequency (NTF), which represents the relative occurrence of the word in the document $d$, as follows:
@@
where $f_{t,d}$ denotes the frequency of word $t$ in document $d$. The $\sum_{t^{\prime}\in d}f_{t^{\prime},d}$ accounts for all words present in $d$. The $N$ represents the total number of occurrences of the word across all documents.
@@
Notably, LM-D and ConDA outperform the others in both accuracy and F1-score. ConDA achieves an accuracy of $0.972$, while the LM-D performs even better, with an accuracy of $0.979$ and an F1-score of $0.980$, making it the most effective detector. Based on these benchmark results, we consider LM-D as the most effective detection method and name LM-D fine-tuned on AIGTBench as OSM-Det, which is subsequently used to quantify and monitor the AAR in social media dataset (SM-D).
@@
Platform # text (Human) FPR
Medium $116,303$ $1.82\%$
Quora $101,145$ $1.36\%$
Reddit $53,321$ $1.70\%$
Table 3: FPR of OSM-Det on social media platforms.
As shown in Table 3, OSM-Det achieves False Positive Rates (FPR) of $1.82\%$, $1.36\%$, and $1.70\%$ on Medium, Quora, and Reddit, respectively, while achieving a benchmark F1-score of $0.980$ (see Table 2). These results highlight OSM-Det’s low misclassification rate and high overall accuracy, making it a reliable choice for quantifying and monitoring AIGTs on social media.
(a) AAR and FPR Trends on Medium from January 1, 2022, to October 31, 2024.
(b) AAR and FPR Trends on Quora from January 1, 2022, to October 31, 2024.
(c) AAR and FPR Trends on Reddit from January 1, 2022, to July 31, 2024.
@@
(c) Word frequency trends on Reddit from January 1, 2022, to July 31, 2024.
Figure 3: Comparison of Medium, Quora, and Reddit word frequency trends: human vs. AI preferences.
@@
We explore the interpretability of the OSM-Det model in the case study using two methods: Integrated Gradients [42], representing a model-dependent perspective, and Shapley Value [36], offering a model-independent perspective. Details of the two methods can be found in Section G.2.
@@
Word-Level Analysis. In the case study of Reddit (refer to  Figures A4 and A6), words like “and”, “think” and “I” have the highest Integrated Gradients and Shapley Value scores, which lead model to classify texts as human-written. Meanwhile, model-specific analysis shows the words “think”, “can”, and “Online” have the lowest scores, leading to AI-generated prediction. From these observations, we note that specifying clear word-level patterns between HWTs and AIGTs is challenging because certain words, like “think”, contribute significantly to both classifications. This overlap suggests that word importance is highly context-dependent, complicating the task of isolating patterns that consistently distinguish the two text types. Similar challenges are also observed on Medium and Quora (refer to Figures A10, A12, A9 and A7).
@@
Given this difficulty, we then turn to a different approach: a statistical analysis of high-frequency adjectives, conjunctions, and adverbs (details provided in Section G.1). These high-frequency terms are then classified into human-preferred and AI-preferred vocabularies. We then track the trends of these lexical items on SM-D.
@@
As shown in Figures 3(a) and 3(b), the NTF of AI-preferred vocabulary on the Medium and Quora is closely aligned with the development of LLMs. Following the release of LLMs such as GPT, Llama, and the Claude series, the NTF of human-preferred vocabulary has gradually declined. Meanwhile, AI-preferred vocabulary shows an increase. These results reflect an increasing usage of LLMs for content generation by Medium and Quora platform users. In contrast, the trends on Reddit show some differences (see Figure 3(c)). From 2022 to 2024, the NTF of human-preferred vocabulary always remains high, while the AI-preferred vocabulary consistently remains low. This indicates that Reddit users rely less on LLMs to produce content.
From above, word frequency changes closely align with the AAR trends shown in Section 5.2.
@@
Sentence-Level Analysis. We also conduct a sentence-level analysis using Shapley values, as Integrated Gradients are only suitable for word-level. From the case studies of Medium, Quora, and Reddit (shown in Figures A11, A8 and A5), we observe that AIGTs are characterized by their objective and standardized structures, typically beginning with a noun or pronoun and following a verb-object pattern, like “Online bullying…contributes…feelings…”. In contrast, HWTs often contain flexible sentence structures and informal expressions, as illustrated by “That being said, why not both?” and “Why can’t we restore…”. In summary, the results suggest that sentence-level patterns provide more distinctive characteristics for distinguishing AIGTs and HWTs, as LLMs may usually follow a standardized pattern to generate texts.
@@
Topic Analysis. Classifying topics on platforms like Quora and Reddit is challenging due to their wide range. Therefore, we focus our analysis on $9$ major topics listed on the Medium platform, [^1] https://medium.com/explore-topics. examining them from a temporal perspective. The proportion of topics is shown in Figure A2.
@@
(a) Number of Likes.
(b) Number of Comments.
Figure 5: Differences between predicted AIGTs and predicted HWTs compressed using a log10 transformation.
@@
As shown in Figure 5(a), the predicted-AIGTs receive fewer “Likes” on average than predicted-HWTs, with mean values of $69.15$ and $127.59$, respectively. And predicted-AIGTs exhibit a higher frequency of low “Likes” counts. Figure 5(b) shows that predicted-AIGTs receive fewer “Comments” on average compared to predicted-HWTs, with mean values of $4.16$ and $7.38$, respectively.

### v2

@@ Abstract
Social media platforms are experiencing a growing presence of AI-Generated Texts (AIGTs). However, the misuse of AIGTs could have profound implications for public opinion, such as spreading misinformation and manipulating narratives. Despite its importance, it remains unclear how prevalent AIGTs are on social media. To address this gap, this paper aims to quantify and monitor the AIGTs on online social media platforms. We first collect a dataset (SM-D) with around $2.4M$ posts from $3$ major social media platforms: Medium, Quora, and Reddit. Then, we construct a diverse dataset (AIGTBench) to train and evaluate AIGT detectors. AIGTBench combines popular open-source datasets and our AIGT datasets generated from social media texts by $12$ LLMs, serving as a benchmark for evaluating mainstream detectors. With this setup, we identify the best-performing detector (OSM-Det). We then apply OSM-Det to SM-D to track AIGTs across social media platforms from January 2022 to October 2024, using the AI Attribution Rate (AAR) as the metric. Specifically, Medium and Quora exhibit marked increases in AAR, rising from $1.77\%$ to $37.03\%$ and $2.06\%$ to $38.95\%$, respectively. In contrast, Reddit shows slower growth, with AAR increasing from $1.31\%$ to $2.45\%$ over the same period. Our further analysis indicates that AIGTs on social media differ from human-written texts across several dimensions, including linguistic patterns, topic distributions, engagement levels, and the follower distribution of authors. We envision our analysis and findings on AIGTs in social media can shed light on future research in this domain.
@@ Introduction
Figure 1: Pipeline of quantifying AIGTs on social media.
@@
Currently, numerous detectors have been developed to detect AIGTs. According to the MGTBench [17], these detectors are broadly divided into two categories: metric-based [12, 36] and model-based detectors [20, 51, 4], some of which have shown high accuracy and robustness. While these detectors have been applied in controlled settings, recent studies have explored their effectiveness in real-world scenarios. Hanley and Durumeric [16] conduct AIGT detection on news website articles, with a primary focus on content generated by GPT-3.5 and others from Turing benchmark, which includes various pre-2022 models [58]. Furthermore, Liu et al. [29] detects ChatGPT-generated content on arXiv papers. However, academic and news writing are formal and tailored to specific audiences, whereas social media content is more interactive, making it a better domain for observing AIGTs’ impact on daily life. Moreover, previous studies do not account for recent popular LLMs, while we consider a broader range of models in our efforts to detect AIGTs on social media.
@@
To quantify and monitor AIGTs on social media, we collect textual data from $3$ popular platforms spanning January 1, 2022, to October 31, 2024, as most LLMs are released after 2022. After data preprocessing, we obtain $1,170,821$ posts from Medium, $245,131$ answers from Quora, and $982,440$ comments from Reddit. We name it as SM-D, short for Social Media Dataset.
@@
To identify the most effective detector, we construct a dataset named AIGTBench, which consists of public AIGT/Supervised-Finetuning (SFT) datasets and our own AIGT datasets generated from social media data. AIGTBench includes AIGTs generated by $12$ different LLMs, such as GPT Series [41]) and Llama Series [56, 57, 10]), totaling around $28.77M$ AIGT and $13.55M$ HWT samples. We then benchmark AIGT detectors on AIGTBench and leverage the best-performing detector as our primary detector, which achieves an accuracy of $0.979$ and an F1-score of $0.980$. To better reflect its application in detecting AIGTs on online social media, we rename it as OSM-Det (Online Social Media Detector).
@@
Based on OSM-Det, we quantify and monitor the texts across the $3$ platforms and use the AI Attribution Rate (AAR) to represent the rate of posts classified as AI-generated (The pipeline is shown in Figure 1). We observe several noteworthy phenomena: (1) A sharp rise in AI-generated content begins in December 2022, with distinct AAR trends emerging across platforms. Before December 2022, the AAR across platforms remains stable. However, starting in December, Medium and Quora show significant surges, while Reddit shows only a slight increase. This suggests the widespread and diverse LLM adoption on social media; (2) Linguistic analysis shows similar AAR trends and exhibits stylistic features in AIGTs/HWTs. Based on the word-level analysis, we find that the usage trend of top-frequency AI-preferred words aligns closely with LLM adoption trends. With sentence-level analysis, we also reveal that AIGTs tend to be more objective and standardized, whereas HWTs are more flexible and informal; (3) Technology-related topics drive higher AARs on Medium. Topics like “Technology” and “Software Development” show the highest AARs, indicating that users with a strong technical background are more likely to adopt LLMs; (4) Predicted HWTs receive more engagement than AIGTs. On Medium, the content predicted as HWTs receives more average “Likes” and “Comments” than AIGTs. This suggests that users are more inclined to engage with HWTs; and (5) Authors with fewer followers are more likely to produce AIGTs. On Medium, users with no more than one thousand followers tend to produce content that has the highest mean AAR at $54.02\%$. In contrast, as the follower count increases, the AAR gradually shifts toward the lower range ($\leq 25.00\%$).
@@
The growth in model parameters and training data has recently empowered LLMs to demonstrate exceptional language processing capabilities [66]. Since then, LLMs have gradually gained popularity, like GPT-4 [40] and Llama [56], enabling users to generate high-quality texts effortlessly. Yet, LLMs have raised concerns about potential misuse, such as fake news generation [65], academic misconduct [59], and performance degradation of training LLMs using AI content [6], making the detection of AIGTs (also known as machine-generated texts) increasingly important [11]. He et al. [17] introduce MGTBench for standardizing the evaluation of different LLMs and experimental setups within the AIGT detectors. They broadly categorize the detectors into two main types: metric-based and model-based detectors. Metric-based detectors use pre-defined metrics, such as log-likelihood, to capture the characteristics of texts [12, 36, 53]. In contrast, model-based detectors rely on trained models to distinguish between AIGTs and HWTs [51, 15, 4, 29, 20, 24]. More introduction refer to Appendix B.
Besides, some researchers have applied detectors to text detection in real-world scenarios. Hanley and Durumeric [16] train a detector using data generated by the ChatGPT and Turing benchmark model and conduct tests on multiple news websites. Their study reveals that, from January 1, 2022, to May 1, 2023, the proportion of synthetic articles increased on news sites. Liu et al. [29] also conduct detection on arXiv and find a significant rise in the proportion of papers using ChatGPT-generated content, reaching $26.1\%$ by December 2023. In contrast to their detection targets, we focus on detecting AIGTs on social media platforms and covering a broader range of LLMs. Macko et al. [30] construct a multilingual dataset based on instant messaging and social interaction platforms such as Telegram and Discord, using it to compare the performance of existing detectors. In contrast, our research focuses on providing an in-depth temporal analysis of AIGTs on content-driven social platforms like Medium, Quora, and Reddit.
@@
Table 1: Overview of the Medium, Quora, and Reddit datasets.
@@
To train the AIGT detectors, we consider two parts of the data. First, we consider $6$ publicly available AIGT datasets and $5$ common SFT datasets to form the training dataset (see Tables A3 and A4 for dataset statistics and Appendix D for more details). Second, to increase the detector’s generalization capabilities on social media, we additionally collect data from the $3$ social media platforms ranging from January 1, 2018, to December 31, 2021. We classify this data as HWTs, given that most LLMs had not been published during this period. We also design different LLMs writing tasks to generate AIGTs that align with the characteristics of platforms (Table A1 describes the statistics details).
@@
Overall, the datasets used for training our detector and the distribution of LLM series are shown in Figure A1. This dataset includes $12$ different LLMs, with a detailed introduction provided in Appendix A. Within these datasets, the two most prevalent model series are the GPT Series, which accounts for $42.99\%$, and the Llama series, which represents $39.05\%$. GPT Series is the most widely used proprietary model and has played a pivotal role in the evolution of generative AI. As of January 2023, approximately $13M$ users interact daily with GPT-3.5 [62]. The Llama series models also have significant influences, as the report indicates that downloads of Llama models on the Hugging Face platform have nearly reached around $350M$ [35]. Therefore, these two model series are the primary focus of our dataset. During the data generation process, we notice that certain samples contain textual noise, like irrelevant or redundant information. To maintain data quality, we implement some data processing strategies (see Appendix E for details).
@@
Table 2: Performance of detectors on AIGTBench. The F1-score corresponds to the AI class.
Experimental Settings
Datasets
As mentioned in Section 3, we collect the social media dataset (SM-D) and the detector training dataset (AIGTBench). SM-D refers to the social media dataset that we conduct the quantification, with more details provided in Section 3.1. AIGTBench is the benchmark for AIGT detectors, which includes AIGTs generated by $12$ different LLMs, as described in Section 3.2. We randomly divide AIGTBench into training, validation, and test sets in a $7:1:2$ ratio. Specifically, the distribution of token lengths in the training, validation, and test set are shown in Figure A2.
AIGT Detectors
During the detection process, we initially use the GPT-2 medium to extract multiple metrics, including log-likelihood and log-rank. Based on these extracted metrics, we train logistic regression models to enhance the accuracy of predictions. For the model-based detectors, we consider both pre-trained detectors and fine-tuned models with the AIGTBench, that is, OpenAI Detector [51], ChatGPT Detector [15], ConDA [4], GPTZero [13], CheckGPT [29], and LM-D [20]. Specifically, for the OpenAI Detector and ChatGPT Detector, we consider their pre-trained version and select the RoBERTa-base model as it demonstrates stable performance across multiple detection tasks and typically provides better detection results. For ConDA and LM-D, we choose the Longformer-base-4096 model as the base model and fine-tune it with the AIGTBench. All of them have a learning rate of 1e-5, a batch size of 16, and the AdamW optimizer. For GPTZero, we directly use its commercial API. For CheckGPT, we retrain the original training framework [29].
@@
We use accuracy and F1-score as the evaluation metrics to evaluate the performance of different detectors, which are common standards in AIGT detection tasks. Besides, we introduce two metrics AI Attribution Rate (AAR) and False Positive Rate (FPR) for quantification analysis. The AAR indicates the proportion of texts that the model predicts as AI-generated, while the FPR denotes the proportion of HWTs misclassified as AIGTs.
To assess word usage, we compute the normalized term frequency (NTF) as:
@@
where $f_{t,d}$ is the frequency of word $t$ in document $d$, $\sum_{t^{\prime}\in d}f_{t^{\prime},d}$ accounts for all words in $d$, and $N$ is the total occurrences of $t$ across all documents.
@@
Notably, LM-D and ConDA outperform the others. ConDA achieves an accuracy of $0.972$, while the LM-D performs even better, with an accuracy of $0.979$ and an F1-score of $0.980$, making it the most effective detector. Based on these benchmark results, we consider LM-D as the most effective detection method and name LM-D fine-tuned on AIGTBench as OSM-Det, which is subsequently used to quantify and monitor the AAR in social media dataset (SM-D). More details on performance across different platforms and all text lengths in SM-D are shown in Appendix G.
Generalizability Experiments. To validate the generalizability of OSM-Det, we conduct experiments under various generation parameter settings, unseen LLMs, and unseen domains. Figure 2 shows that OSM-Det can maintain excellent performance on AIGTs under different generation parameters (temperature, Top-P and Top-K), with accuracy exceeding $0.990$. The generalization performance on unseen LLMs and domains stays robust. For further details, please refer to Appendix H. Thus, these results show that OSM-Det has outstanding generalizability, supporting the reliability of the results on SM-D.
Figure 2: Impact of different generation parameters on AIGT detection accuracy.
@@
As shown in Table A2, OSM-Det achieves False Positive Rates (FPR) of $1.82\%$, $1.36\%$, and $1.70\%$ on Medium, Quora, and Reddit, respectively, while achieving a benchmark F1-score of $0.980$ (see Table 2). These results highlight OSM-Det’s low misclassification rate and high overall accuracy, making it a reliable choice for quantifying and monitoring AIGTs on social media.
(a) AAR Trends on Medium from January 1, 2022, to October 31, 2024.
(b) AAR Trends on Quora from January 1, 2022, to October 31, 2024.
(c) AAR Trends on Reddit from January 1, 2022, to July 31, 2024.
@@
Figure 4: Comparison of Medium and Quora word frequency trends: human vs. AI preferences. (The result of Reddit is shown in Figure A5)
@@
We explore the interpretability of the OSM-Det model in the case study using two methods: Integrated Gradients [54], representing a model-dependent perspective, and Shapley Value [48], offering a model-independent perspective. Details of the two methods can be found in Section I.2.
@@
Word-Level Analysis. In the case study of Reddit (refer to  Figures A8 and A10), words like “and”, “think” and “I” have the highest Integrated Gradients and Shapley Values, which lead model to classify texts as human-written. Meanwhile, model-specific analysis shows the words “think”, “can”, and “Online” have the lowest scores, leading to AI-generated prediction. From these observations, we note that specifying clear word-level patterns between two class is challenging because certain words, like “think”, contribute significantly to both classifications. This overlap suggests that word importance is highly context-dependent. Similar challenges are also observed on Medium and Quora (Figures A14, A16, A13 and A11).
@@
Given this difficulty, we then turn to a different approach: a statistical analysis of high-frequency adjectives, conjunctions, and adverbs (details provided in Section I.1). These high-frequency terms are then classified into human-preferred and AI-preferred vocabularies. We then track the trends of these lexical items on SM-D.
@@
As shown in Figures 4(a) and 4(b), the NTF of AI-preferred vocabulary on the Medium and Quora is closely aligned with the development of LLMs. Following the release of LLMs such as GPT, Llama, and the Claude series, the NTF of human-preferred vocabulary has gradually declined. Meanwhile, AI-preferred vocabulary shows an increase. These results reflect an increasing usage of LLMs for content generation by Medium and Quora platform users. In contrast, the trends on Reddit show some differences (see Figure A5). From 2022 to 2024, the NTF of human-preferred vocabulary always remains high, while the AI-preferred vocabulary consistently remains low. This indicates that Reddit users rely less on LLMs to produce content. From above, we obverse that word frequency changes closely align with the AAR trends in Figure 3.
@@
Sentence-Level Analysis. We also conduct a sentence-level analysis using Shapley values, as Integrated Gradients are only suitable for word-level. From the case studies of Medium, Quora, and Reddit (shown in Figures A15, A12 and A9), we observe that AIGTs are characterized by their objective and standardized structures, typically beginning with a noun or pronoun and following a verb-object pattern, like “Online bullying…contributes…feelings…”. In contrast, HWTs often contain flexible sentence structures and informal expressions, as illustrated by “That being said, why not both?” and “Why can’t we restore…”. In summary, the results suggest that sentence-level patterns provide more distinctive characteristics for distinguishing AIGTs and HWTs, as LLMs may usually follow a standardized pattern to generate texts.
@@
Topic Analysis. Classifying topics on platforms like Quora and Reddit is challenging due to their wide range. Therefore, we focus our analysis on $9$ major topics listed on the Medium [33], examining them from a temporal perspective. The proportion of topics is shown in Figure A3.
@@
Follower Group Mean Likes (AIGTs / HWTs) Mean Comments (AIGTs / HWTs)
0-1K $49.48/79.39$ $3.18/5.68$
1-5K $111.50/191.61$ $5.11/9.09$
>5K $126.94/211.92$ $5.56/8.25$
Table 3: Engagement statistics on Medium for different follower groups, comparing AIGTs and HWTs.
@@
As shown in Figure 4(a), the predicted-AIGTs receive fewer “Likes” on average than predicted-HWTs, with mean values of $69.15$ and $127.59$, respectively. And predicted-AIGTs exhibit a higher frequency of low “Likes” counts. Figure 4(b) shows that predicted-AIGTs receive fewer “Comments” on average compared to predicted-HWTs, with mean values of $4.16$ and $7.38$, respectively. We further investigate the mean values of Likes and Comments for authors with different numbers of followers and Table 3 indicates that, across all follower count groups, AIGTs receive significantly fewer Likes and Comments compared to HWTs.